# A theoretical analysis on the resolution of parametric PDEs via neural networks designed with Strassen algorithm

### v1

@@ 2.2 Strassen Algorithm for Square Matrices
In this section we recall the Strassen algorithm, which was proposed in [Str69]. First of all, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
Let us now define recursively the Strassen algorithm for matrices in $\mathbb{R}^{2^{k}\times 2^{k}}$. The base case $k=0$, simply consists on multiplications of scalars.
Let us consider $k\geq 1$ and suppose that we know how to perform Strassen multiplication for matrices in $\mathbb{R}^{2^{k}\times 2^{k}}$ and obtain it for matrices in $\mathbb{R}^{2^{k+1}\times 2^{k+1}}$. Then, we obtain that:
@@
$\begin{split}P_{1}&=(A[1,1]+A[2,2])(B[1,1]+B[2,2]),\\ P_{2}&=(A[2,1]+A[2,2])B[1,1],\\ P_{3}&=A[1,1](B[1,2]-B[2,2]),\\ P_{4}&=A[2,2](B[2,1]-B[1,1]),\\ P_{5}&=(A[1,1]+A[1,2])B[2,2],\\ P_{6}&=(A[2,1]-A[1,1])(B[1,1]+B[1,2]),\\ P_{7}&=(A[1,2]-A[2,2])(B[2,1]+B[2,2]).\end{split}$ (2.2)
@@
    It is then enough to apply Strassen to this extended matrix and recover the convenient submatrix.
@@
    We consider a division of blocks of size $2^{k}\times 2^{k}$, perform the multiplication with such blocks, and do the corresponding additions. For example, if $n=6$ and $A,B\in\mathbb{R}^{6\times 6}$ then we can express the product $AB$ as a sum of products of submatrices where we can apply the Strassen algorithm:
    $(AB)_{1:4,1:4}$ $=A_{1:4,1:4}B_{1:4,1:4}+\begin{pmatrix}A_{1:2,5:6}B_{5:6,1:2}&A_{1:2,5:6}B_{5:6,3:4}\\ A_{3:4,5:6}B_{5:6,1:2}&A_{1:2,5:6}B_{5:6,3:4}\end{pmatrix},$
@@ 2.3 Matrix Neural Networks
We now define the notions of matrix neural network, concatenation, and parallelization. This is based on the definition introduced in [PV18], but adapted for neural networks whose entries are matrices.
@@
where $N^{1,1},N^{1,2},\dots,N^{L,1},N^{L,2}\in\mathbb{N}$, and where each $\mathcal{L}^{\ell}$ is a $(N^{\ell,1}\times N^{\ell,2})\times(N^{\ell-1,1}\times N^{\ell-1,2})$ tensor, $C^{\ell}$ is a $N^{\ell,1}\times N^{\ell,2}$ matrix and
@@
which we call the number of weights of $\Phi$. Moreover, we refer to $\operatorname{dim_{out}}\Phi\coloneq N^{L,1}\times N^{L,2}$ as the output dimension of $\Phi$.
Given a previously defined Matrix Neural Network $\Phi$, in order to work with its layers, we write $((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},B^{L},\alpha^{L}))\coloneq\Phi$ to specify which and in what order tensors, matrices, and activation functions conform to the matrix neural network.
A classical Neural Network has the same structure as a Matrix Neural Network: it is a sequence of triples of matrices, vectors, and activation functions. Its realization, number of layers and number of weights are defined equivalently for these Neural Networks.
@@
Any classical neural network (if we consider them defined as in [PV18]), can be viewed as a matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Moreover, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the matrix neural networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent notions.
@@
such that $\operatorname{dim_{in}}\Phi^{1}=\operatorname{dim_{out}}\Phi^{2}$, the $\varrho$-Matrix Neural Network $\phi^{1}\bullet\phi^{2}$ is defined as
@@
This $\varrho$-MNN satisfies again the property $R\left(\Phi^{1}\bullet\Phi^{2}\right)=R\left(\Phi^{1}\right)\circ R\left(\Phi^{2}\right)$. Even though this definition of concatenation produces Matrix Neural Networks with fewer layers, the estimation on the number of weights generally gets worse. This leads us to not use this definition and, therefore, not introduce it.
Note that if $\Phi^{1}$ and $\Phi^{2}$ are a $\varrho$-MNN then $\Phi^{1}\circ\Phi^{2}$ is also a $\varrho$-MNN too.
@@
Animportant example is $\varrho(t)=\max\left\{0,t\right\}^{2}$. In this case, the product is a $\varrho$-realization:
@@
Thus, for $\varrho(t)=\max\left\{0,t\right\}^{2}$ there is a neural network satisfying:
@@
and computes the product exactly. This network is

### v2

@@ 2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [Str69]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
Let us now recursively define the Strassen algorithm for the matrices in $\mathbb{R}^{2^{k}\times 2^{k}}$. The base case $k=0$ simply consists of multiplications of scalars. The inductive case, $k\geq 1$, is defined as follows:
@@
$\begin{split}P_{1}&=(A[1,1]+A[2,2])(B[1,1]+B[2,2]),\\ P_{2}&=(A[2,1]+A[2,2])B[1,1],\\ P_{3}&=A[1,1](B[1,2]-B[2,2]),\\ P_{4}&=A[2,2](B[2,1]-B[1,1]),\\ P_{5}&=(A[1,1]+A[1,2])B[2,2],\\ P_{6}&=(A[2,1]-A[1,1])(B[1,1]+B[1,2]),\\ P_{7}&=(A[1,2]-A[2,2])(B[2,1]+B[2,2]),\end{split}$ (2.3)
where all the multiplications in (2.3) consist of the Strassen multiplication in $\mathbb{R}^{2^{k-1}\times 2^{k-1}}$.
@@
    It is then enough to apply the Strassen multiplication to this extended matrix and recover the convenient submatrix.
@@
    We consider a division of blocks of size $2^{k}\times 2^{k}$, perform the multiplication with such blocks, and do the corresponding additions. For example, if $n=6$ and $A,B\in\mathbb{R}^{6\times 6}$, we can express the product $AB$ as a sum of products of submatrices where we can apply the Strassen algorithm:
    $(AB)_{1:4,1:4}$ $=A_{1:4,1:4}B_{1:4,1:4}+\begin{pmatrix}A_{1:2,5:6}B_{5:6,1:2}&A_{1:2,5:6}B_{5:6,3:4}\\ A_{3:4,5:6}B_{5:6,1:2}&A_{3:4,5:6}B_{5:6,3:4}\end{pmatrix},$
@@ 2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [PV18], but adapted for Neural Networks whose entries are matrices. We do that for the sake of simplifying the notation.
@@
where $N^{1,1},N^{1,2},\dots,N^{L,1},N^{L,2}\in\mathbb{N}$, and where each $\mathcal{L}^{\ell}$ is a
$(N^{\ell,1}\times N^{\ell,2})\times(N^{\ell-1,1}\times N^{\ell-1,2})$
tensor, $C^{\ell}$ is a $N^{\ell,1}\times N^{\ell,2}$ matrix and
@@
which we call the number of weights of $\Phi$. Moreover, we refer to
$\operatorname{dim_{out}}\Phi\coloneq N^{L,1}\times N^{L,2}$
as the output dimension of $\Phi$.
Finally, given a previously defined Matrix Neural Network $\Phi$, in order to work with its layers, we write $((\mathcal{L}^{1},B^{1},\alpha^{1}),\dots,(\mathcal{L}^{L},B^{L},\alpha^{L}))\coloneq\Phi$ to specify which and in what order tensors, matrices, and activation functions conform to the Matrix Neural Network.
@@
Classical Neural Networks (if we consider them defined as in [PV18]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any classical Neural Network can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
@@
such that $\operatorname{dim_{in}}\Phi^{1}=\operatorname{dim_{out}}\Phi^{2}$, the $\varrho$-Matrix Neural Network $\Phi^{1}\bullet\Phi^{2}$ is defined as:
@@
This $\varrho$-MNN satisfies again the property $R\left(\Phi^{1}\bullet\Phi^{2}\right)=R\left(\Phi^{1}\right)\circ R\left(\Phi^{2}\right)$. Even though this definition of concatenation produces Matrix Neural Networks with fewer layers, the estimation on the number of weights generally gets worse. This leads us to not use this definition.
Note that if $\Phi^{1}$ and $\Phi^{2}$ are a $\varrho$-MNN, then $\Phi^{1}\circ\Phi^{2}$ is also a $\varrho$-MNN too.
@@
An important example is ReLU, as we show in Section 6.
Another relevant example is $\varrho(t)=\max\left\{0,t\right\}^{2}$. In this case, the product is a $\varrho$-realization:
@@
Thus, for $\varrho(t)=\max\left\{0,t\right\}^{2}$ there is a Neural Network $\Pi^{\varrho}_{0,\infty}$ satisfying:
@@
and computes the product exactly. This network is:
@@ 3.1 Strassen Algorithm for Square Matrices Whose Size is a Power of 2 Using Neural Networks
$R\left(\Phi^{\operatorname{STR}_{2^{0},\varrho}}_{\varepsilon,K}\right)(A|B)=R\left(\Pi_{\varepsilon,K}^{\varrho}\right)(A_{1,1},B_{1,1}),$ (3.1)
for $\Pi_{\varepsilon,K}^{\varrho}$ defined in Hypothesis 1.
For the recursive case, we need to define several auxiliary networks. The first, whose input is in $\mathbb{R}^{(7\cdot 2^{k-1})\times 2^{k-1}}$ and whose output is in $\mathbb{R}^{2^{k}\times 2^{k}}$, is given by:
$R\left(\Phi^{\operatorname{MIX}_{k}}\right)\begin{bmatrix}P_{1}\\ P_{2}\\ P_{3}\\ P_{4}\\ P_{5}\\ P_{6}\\ P_{7}\end{bmatrix}=\begin{pmatrix}P_{1}+P_{4}-P_{5}+P_{7}&P_{3}+P_{5}\\ P_{2}+P_{4}&P_{1}-P_{2}+P_{3}+P_{6}\end{pmatrix}.$
Note that $\Phi^{\operatorname{MIX}_{k}}$ is a one layer MNN such that:
$M\left(\Phi^{\operatorname{MIX}_{k}}\right)=12\cdot(2^{k-1})^{2}=3\cdot 4^{k}.$ (3.2)
Also, with the notation of Definition 2.2, we define a MNN whose input is in $\mathbb{R}^{2^{k}\times 2^{k+1}}$ and whose output is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2\cdot 2^{k-1}}$:
$R(\Phi^{\operatorname{SPLIT}_{k}})(A|B)=\begin{pmatrix}A[1,1]+A[2,2]&|&B[1,1]+B[2,2]\\ A[2,1]+A[2,2]&|&B[1,1]\\ A[1,1]&|&B[1,2]-B[2,2]\\ A[2,2]&|&B[2,1]-B[1,1]\\ A[1,1]+A[1,2]&|&B[2,2]\\ A[2,1]-A[1,1]&|&B[1,1]+B[1,2]\\ A[1,2]-A[2,2]&|&B[2,1]+B[2,2]\end{pmatrix}.$
This network satisfies:
$M(\Phi^{\operatorname{SPLIT}_{k}})=24\cdot(2^{k-1})^{2}=6\cdot 4^{k}.$ (3.3)
Finally, we define a MNN whose input is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2\cdot 2^{k-1}}$ and whose output is in $\mathbb{R}^{7\cdot 2^{k-1}\times 2^{k-1}}$:
$\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\right.\\ \left.\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K}\right).$
Bearing this in mind, we define recursively that:
$\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}=\Phi^{\operatorname{MIX}_{k}}\circ\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}\circ\Phi^{\operatorname{SPLIT}_{k}}.$
Let us show the following result with respect to the accuracy and size of the MNN that we have designed:
Proposition 3.1.
Let $k\in\mathbb{N}\cup\left\{0\right\}$, $A,B\in\mathbb{R}^{2^{k}\times 2^{k}}$, $\varepsilon>0$, and $K>0$ satisfying:
$\left\lvert A\right\rvert_{\infty},\left\lvert B\right\rvert_{\infty}\leq K.$
Then,
$\left\lvert R\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)(A|B)-AB\right\rvert_{\infty}\leq\varepsilon,$ (3.4)
$M\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)=7^{k}\left(M\left(\Pi^{\varrho}_{\frac{\varepsilon}{4^{k}},2^{k}K}\right)+12\right)-12\cdot 4^{k},$ (3.5)
$L\left(\Phi^{\operatorname{STR}_{2^{k},\varrho}}_{\varepsilon,K}\right)=L\left(\Pi_{\frac{\varepsilon}{4^{k}},2^{k}K}^{\varrho}\right)+2k.$ (3.6)
Proof.
The statement (3.4) can be proved inductively on $k$. The base case $k=0$, when $A$ and $B$ are $1\times 1$ matrices, that is, scalars, is a direct consequence of (3.1) and Hypothesis 1.
Let us suppose that Proposition 3.1 is true for $k-1$. Let us start by showing (3.4). Using the Strassen algorithm, we know that (2.2) is satisfied. Consequently, because of the triangular inequality, defining $P_{i}$ as in (2.3), it suffices to show that: